OpenStudy (anonymous):

MEDAL REWARDED Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 4, -2, and -1 + 2i

OpenStudy (anonymous):

@Hero if you would like to help.

hero (hero):

Is this a question you've answered already?

OpenStudy (anonymous):

I have, but im not sure im right, i get confused in some of the steps sometimes

hero (hero):

Okay, so what answer did you get for this?

OpenStudy (anonymous):

f(x) = x4 - 3x3 + 8x2 + 13x + 40

hero (hero):

So here's what you're supposed to do for these: They give you the zeroes so you do this: x = -2 Then add 2 to both sides: x + 2 = 0 x = 4 Subtract 4 from both sides x - 4 = 0 x = -1 - 2i x = -1 + 2i then add the terms on the left to the other side to get x + 1 + 2i = 0 x + 1 - 2i = 0 Then you take all the terms that have been set equal to zero, and multiply them (x + 2)(x - 4)(x + 1 + 2i)(x + 1 - 2i) After doing so, you get: \(x^4-7 x^2-26 x-40\)

OpenStudy (anonymous):

I did some multiplying wrong...

hero (hero):

Are you sure you multiplied the right factors?

OpenStudy (anonymous):

Thank you though, i see what i did wrong. Every question i post, i have multiply that are the same with different numbers so after i understand it i go and do the other ones on my own.